OpenStudy (anonymous):

r(x)=(x+1)/(x(x+4)) what is the end behavior near -1???

OpenStudy (kira_yamato):

It approaches the value 0

OpenStudy (anonymous):

how did you get that?

OpenStudy (kira_yamato):

Plug in x = -1, and you get r(-1) = (-1+1)/(-1(-1+4)) = 0/(-3) = 0 Hence as x approaches -1, r(x) approaches 0

OpenStudy (anonymous):

Are we talking about limits? I guess he takes the limits as x goes to -1. Plug the -1 into the equations then u get zero

OpenStudy (anonymous):

i hope you can help me figure this out. in the book it shows this r(x)≈ -1/3(x+1) after plugging in the -1. Any idea how they got that?

OpenStudy (kira_yamato):

No idea....